OpenStudy (anonymous):

two point charges of +5.0 * 10^-5 C and -5.0 * 10^-5 C are separated by 0.50 m. Which of the following describes the force between them? A. 90 N , attractive B. 45 N, repulsive C. 45 N , attractive D. 90 N , repulsive

OpenStudy (michele_laino):

that force is attractive, since the charge have opposite sign

OpenStudy (anonymous):

yes:) how can we find if it is 45 or 90?

OpenStudy (michele_laino):

furthermore, we have to apply the Coulomb law: \[\Large F = K\frac{{{Q_1}{Q_2}}}{{{r^2}}} = 9 \times {10^9}\frac{{5 \times {{10}^{ - 5}} \times 5 \times {{10}^{ - 5}}}}{{{{\left( {0.5} \right)}^2}}}\]

OpenStudy (anonymous):

so we get 90? so option A is our solutoin?

OpenStudy (michele_laino):

yes! correct!

OpenStudy (anonymous):

yay!:)thanks!! okay onto the next! :P

OpenStudy (michele_laino):

:)